1. Plains States Manufacturing has just signed a contract to sell agricultural equipment to Boschin, a German firm, for €1,250,000. The sale was made in June with payment due six months later in December. Because this is a sizable contract for the firm and because the contract is in Euros rather than dollars, Plains States is considering several hedging alternatives to reduce the exchange rate risk arising from the sale. To help the firm make a hedging decision you have gathered the following information.

The spot exchange rate is $0.8924/€

The six month forward rate is $0.8750/€

Plains States’ cost of capital is 11%

The Euro zone 6-month borrowing rate is 9% (or 4.5% for 6 months)

The Euro zone 6-month lending rate is 7% (or 3.5% for 6 months)

The U.S. 6-month borrowing rate is 8% (or 4% for 6 months)

The U.S. 6-month lending rate is 6% (or 3% for 6 months)

Plains States’ forecast for 6-month spot rates is $0.91/€

The budget rate, or the lowest acceptable sales price for this project, is $1,075,000 or $0.86/€

a. If Plains States chooses not to hedge their euro receivable, the amount they receive in six months will be __________.

b. If Plains States chooses to hedge its transaction exposure in the forward market, it will __________ €1,250,000 forward at a rate of __________. If Plains States chooses to hedge its transaction exposure in the forward market at the available forward rate. How much will it receive in 6 months?

c. Plains States could hedge the Euro receivables in the money market. If it chooses the money market hedge, how much would it receive in six months assuming Plains States reinvests the proceeds at the U.S. investment rate?
Answer #1

a). Without hedging, Plains State will receive amount receivable*6-month spot rate

= €1,250,000*$0.91/€ = $1,137,500 (This is uncertain since exchange rates are subject to change.)

b).  If Plains States chooses to hedge its transaction exposure in the forward market, it will sell €1,250,000 forward at a rate of $0.8750/€. It will receive 1,250,000*0.8750 = $1,093,750.

c). Money market hedge:

6-month borrowing rate in Euro is 4.5% so amount to be borrowed now = €1,250,000/(1+4.5%) = €1,196,172

Loan to be given now = current spot rate*€1,196,172 = 0.8924*1,196,172 = $1,067,424

Reinvestment rate for 6 months = US lending rate = 6%/2 = 3%

Settlement amount in 6 months = $1,067,424*(1+3%) = $1,099,488